OpenStudy (anonymous):

Choose the correct graph of the given system of equations. 3y − x = 4 y + 2x = −1 A) One solution (-1,1) B) One solution (1,1) C) No solution D) None of the above

OpenStudy (radar):

X=-1, y=1 A) looks like a good choice.

OpenStudy (anonymous):

A

OpenStudy (radar):

One solution (-1,1)

OpenStudy (radar):

A) (-1,1)

OpenStudy (anonymous):

Ok thanks guys. :) And one more.. Kinda long.. Part 1: Provide a system of TWO equations in slope-intercept form, with only one solution. Using complete sentences, explain why this system has one solution. Part 2: Provide a system of TWO equations in slope-intercept form with no solutions. Using complete sentences, explain why this system has no solutions. Part 3: Provide a system of TWO equations in slope-intercept form with infinitely many solutions. Using complete sentences, explain why this system has infinitely many solutions.

OpenStudy (anonymous):

You dnt have to do the complete sentances part, thats too much..

OpenStudy (radar):

Which part do you want to do first?

OpenStudy (anonymous):

1 please

OpenStudy (radar):

You do know what the slope intercept form looks like and we need to provide two equations

OpenStudy (anonymous):

mhmm, y=mx+b

OpenStudy (radar):

I am having trouble posting my answers. It is taking a long time. y=3x-16 y=-2x+14 I believe these two only have one solution.

OpenStudy (radar):

The solution is x=6 and y =2. There is only one solution because these two lines only intersect at one point (6,2)

OpenStudy (radar):

That should take care of Part 1:

OpenStudy (radar):

You ready for Part 2:???

OpenStudy (radar):

y=3x-2 y=3x+2 No solution parallel lines that never intersect. Thats Part 2.

OpenStudy (radar):

You take care of Part 3. Bye and good luck.